The demand and supply equations for a certain item are given by
D = –5p + 40 S = –p^2 + 30p – 8
Find the equilibrium price

Answer by stanbon(57377)

The demand and supply equations for a certain item are given by
D = –5p + 40 S = –p^2 + 30p – 8
Find the equilibrium price
-----------------------
-p^2+30p-8 = -5p+40
p^2-35p+48=0
p = [35 +- sqrt(35^2-4*48)]/2
p = [35 +- sqrt(1033)]/2
p = [35 +- 32.14]/2
p = 33.57 or p = 1.43
==========================
Cheers,
Stan H.